OpenStudy (anonymous):

Carl can paint a room 3 hours faster than Jennifer can. If they work together, they can complete the job in 2 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Jennifer to complete this job on her own.

OpenStudy (anonymous):

i really need help asap!!!!

OpenStudy (anonymous):

Ok, so is carl can paint a room 3x the speed that jennifer can, we can label jennifer's speed as x, and they can complete the job in 2 hours. Therefore, set up the equation as \[1/x + 1/(x -3) = 1/2\] Now, multiply both sides by 2 to get \[1/x (2) + 1/(x-3) (2) = 1\] Now, solve for x to find how long it would take for jennifer to paint the room.\[2/x + 2/(x-3) = 1 \] After solving, you would get x =6,1. Therefore, it would take jennifer 6 hours to complete the job on her own

OpenStudy (anonymous):

Is 6 an answer choice?

OpenStudy (anonymous):

thank you @desiswag

OpenStudy (anonymous):

aha yw :) Are you taking flvs?

OpenStudy (anonymous):

yes its hard for me i still have other problems i dont know

OpenStudy (anonymous):

Algebra 2?

OpenStudy (anonymous):

Im taking that course too if you are so I could help you on some other problems :)

OpenStudy (anonymous):

yess and really thank you i have no one to help me so its kind of hard for me and the examples dont help

OpenStudy (anonymous):

When looking at a rational function, Charles and Bobby have two different thoughts. Charles says that the function is defined at x = −2, x = 3, and x = 5. Bobby says that the function is undefined at those x values. Describe a situation where Charles is correct, and describe a situation where Bobby is correct. Is it possible for a situation to exist where they are both correct? Justify your reasoning. do you know this one by any chance@desiswag

OpenStudy (anonymous):

What lesson is this?

OpenStudy (anonymous):

Ok, so all you would have to do for this problem is you would have to come up with a rational function that has those three x values as solutions, and a function that has those three x values as extraneous solutions

OpenStudy (anonymous):

5.08 and how would i do that im confused

OpenStudy (anonymous):

oh lol im on 5.07 coincidence :P http://openstudy.com/study#/updates/52c57e44e4b0b729fb8c9981 try that one mathmale explains things better than I ever could

OpenStudy (anonymous):

thanks are you doing the essay part of the test already?

OpenStudy (anonymous):

Lol no not yet

OpenStudy (anonymous):

i am its pretty hard i still got 2 left and ive been here for hours im so lost @desiswag

OpenStudy (anonymous):

It takes me anywhere from 4 - 7 hours to do the essay portion of the tests I almost go insane :|

OpenStudy (anonymous):

same here and my teacher grades so hard i always have to redo the tests

OpenStudy (anonymous):

Ikr!! My average went down though :(

OpenStudy (anonymous):

really? yeah im about to get a c cause of these tests

OpenStudy (anonymous):

It's really hard for me because my math teacher at school wanted me to take it so I can do pre- calculus in 8th grade next year so I really didn't get a break from school :\

OpenStudy (anonymous):

wow your so smart then you could do it!!!

OpenStudy (anonymous):

could you help me on something again sorry to bother you

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

on the first question i posted how would i keep solving 2/x+2/x-3=1

OpenStudy (anonymous):

so what you would do first is you would find a common denominator. What factors do x and x-3 have in common?

OpenStudy (anonymous):

dont they have in common x?

OpenStudy (anonymous):

x is common to both denominators, so multiply x(x-3). That is the common denominator. Now, just like finding the common denominator of simple fractions,